• An aircraft attempt to fly due north at 100km per hour.if the wind blows against it from east to west at 60km per hour.it resultant velocity is?
1
Expert's answer
2022-01-30T13:51:15-0500


The resultant velocity

"V=\sqrt{100^2+40^2}=108\:\rm km/h"

"\theta=\tan^{-1} \frac{100}{40}=68^\circ\quad \rm N\;of\;W"